OpenStudy (anonymous):

Find the time it takes an objecy to hit the ground when it is dropped from a height of s feet. h=-16t^2+s s=80

OpenStudy (anonymous):

Set h equal to 0 and solve for t.

OpenStudy (campbell_st):

ok.... the subsititute s = 80 and let h = 0 0 = =16t^2 + 80 16t^2 = 80 t^2 = 5 now solve for t to get the time...

OpenStudy (lgbasallote):

when the object hits the ground then h = 0 since it's given s= 80 just substitute 0 = -16t^2 + 80 16t^2 = 80 solve for t

OpenStudy (lgbasallote):

does that help?

OpenStudy (anonymous):

yeah thanks